solve by the substitution method.
2x - 7y=15
5x + 45=y

Answer by jsmallt9(3758)

2x - 7y=15
5x + 45=y

When looking to use the Substitution Method, whether by choice or by direction, you look for an equation where one of the variables is already solved for (or easy to solve for). In your system, the second equation is already solved for y. So we will use that expression for y and substitute it in for y in the first equation:
2x - 7(5x + 45)=15
Note the use of parentheses! This is an extremely good habit to have when substituting one expression for another! In this case it helps us understand that the 7 should be distributed:
2x -35x - 315 = 15
Simplifying the left side:
-33x - 315 = 15
Add 316:
-33x = 330
Divide by -33:
x = -10

This is half of the solution. We also need the y value:
5(-10) + 45 = y
-50 + 45 = y
-5 = y

So the solution is: (-10, -5)